Consider the reaction when aqueous solutions of iron(II) sulfate and lead(II) nitrate are combined. The net ionic equation for this reaction is:
1
Expert's answer
2019-03-05T07:20:59-0500

Answer:

FeSO4 + Pb(NO3)2 = PbSO4 + Fe(NO3)2

Fe^2+ + SO4^2- + Pb^2+ + 2NO3^- = PbSO4 + Fe2+ + 2NO3^-

Pb^2+ + SO4^2- = PbSO4